OpenStudy (anonymous):

If Barbara received a 4 percent raise and is now making $41,200 per year, what was her salary before the raise? Round your answer to the nearest dollar.

OpenStudy (anonymous):

Well let us make a variable we can use x = salary before the raise We don't know what x is yet but we do know it is some number. Let's multiply her raise by 4 percent and make it equal 41, 200 remeber that 4% is 0.04 in decimal form. We are going to multiply by 1.04 because we want her salary plus 4% of it where as by multiplying by just 0.04 we would only get 4% of her salary x * 1.04 = 41,200 Now let's solve for x x * 1.04 =41,200 x * 1.04 *1 /1.04 = 41,200 / 1.04 divide by 1.04 on both sides x = 39,615.3 round x = 39,615 Lets check our answer 39, 615 * 1.04 = 41,199.6

OpenStudy (anonymous):

thanks a lot!!!!!!!!!!!!